OpenStudy (anonymous):

What is the slope of the line passing through the point (3, −4) and perpendicular to the line y = 3x + 5? Write your answer as a fraction.

OpenStudy (anonymous):

so u have point (3,-4) and the gradient of the perpendicular line is gonna be -1/gradient of the line y=3x+5 so -1/3 .. use y-y1=m(x-x1) to find equation of perpendicular line

OpenStudy (anonymous):

so y-(-4)=m(x-3)?

hartnn (hartnn):

isn't product of slopes of two perpendicular lines=-1.....and answer simply be -1/3??

OpenStudy (anonymous):

idk thats why I'm asking.

OpenStudy (anonymous):

anyone know the answer to this???

OpenStudy (anonymous):

oh my bad .. its just asking for slope of the line .. thought it was asking for equation

OpenStudy (anonymous):

no just the slope

OpenStudy (anonymous):

know how to find the slope of this @powerangers69 ?

OpenStudy (anonymous):

um for 2 linear lines to be parallel their gradients m1xm2=-1 so u have m1 just sub that in and solve for m2 which is slope of perpendicular line

OpenStudy (anonymous):

so it would be -1/3?

OpenStudy (anonymous):

yeah